OpenStudy (anonymous):

A ship starting 500 nautical miles from port travels at a speed of 27 knots or less. How long does the trip to port take? HINT: A trip may take more time if the ship travels slower than its top speed of 27 knots.

OpenStudy (anonymous):

Use the Formula: S=vt S=distance , v=Speed and t=time t=S/v t=500/27 t=18.51